# Discrete Probabilistic Inference as Control in Multi-path Environments

### v1

@@ 1 Introduction
Approximate probabilistic inference has seen a tremendous amount of progress, notably from a variational perspective coupled with deep neural networks. This is particularly true for continuous sample spaces, where the Evidence Lower Bound (ELBO) can be maximized with gradient methods, thanks to methods such as pathwise gradient estimation (Kingma and Welling, 2014; Rezende et al., 2014; Rezende and Mohamed, 2015). In the case of discrete and highly structured sample spaces though, this “reparametrization trick” becomes more challenging since it often requires continuous relaxations of discrete distributions (Jang et al., 2017; Maddison et al., 2017; Mena et al., 2018). In those cases, variational inference can also be carried out more generally thanks to score function estimation (Williams, 1992), albeit at the expense of high variance (Mohamed et al., 2020). An alternative for approximate inference is through sampling methods, based on Markov chain Monte Carlo (MCMC; Hastings, 1970; Gelfand and Smith, 1990), when the target distribution is defined up to an intractable normalization constant.
@@
Bengio et al. (2021) introduced a new class of probabilistic models called Generative Flow Networks (GFlowNets), to approximate an unnormalized target distribution over discrete and structured sample spaces from a variational perspective (Malkin et al., 2023; Zimmermann et al., 2022). GFlowNets treat sampling as a sequential decision making problem, heavily inspired by the literature in Reinforcement Learning (RL). Unlike RL though, which seeks an optimal policy maximizing the cumulative reward, the objective of a GFlowNet is to find a policy such that objects can be sampled proportionally to their cumulative reward. Nevertheless, this relationship led to a number of best practices from the RL literature being transferred into GFlowNets, such as the use of a replay buffer (Shen et al., 2023; Vemgal et al., 2023) and target network (Deleu et al., 2022), and advanced exploration strategies (Rector-Brooks et al., 2023).
@@
Although Bengio et al. (2021) proved that GFlowNets were exactly equivalent to maximum entropy RL (MaxEnt RL; Ziebart, 2010) in some specific cases which we will recall in Section 2, it has long been thought that this connection was only superficial in general. However recently, Tiapkin et al. (2024) showed that GFlowNets and MaxEnt RL are in fact one and the same, up to a correction of the reward function. This, along with other recent works (Mohammadpour et al., 2024; Anonymous, 2023), paved the way to show deeper connections between GFlowNets and MaxEnt RL algorithms. In this work, we extend this correction of the reward to a more general case, and establish novel equivalences between GFlowNet & MaxEnt RL objectives, notably between the widely used Trajectory Balance loss in the GFlowNet literature (Malkin et al., 2022), and the Path Consistency Learning algorithm in MaxEnt RL (Nachum et al., 2017). We also introduce a variant of the Soft Q-Learning algorithm (Haarnoja et al., 2017), depending directly on a policy, and show that it becomes equivalent to the Modified Detailed Balance loss introduced by Deleu et al. (2022). Finally, we show these similarities in behavior empirically on three different domains.
@@ 2.1 Maximum entropy Reinforcement Learning
We consider a finite-horizon Markov Decision Process (MDP) ${\mathcal{M}}_{\mathrm{soft}}=({\mathcal{S}},{\mathcal{A}},s_{0},T,r)$, where the state space ${\mathcal{S}}$ and the action space ${\mathcal{A}}$ are discrete and finite. We assume that this MDP is deterministic, with a transition function $T:{\mathcal{S}}\times{\mathcal{A}}\rightarrow{\mathcal{S}}$ that determines how to move to a new state $s^{\prime}=T(s,a)$ from the state $s$, following the action $a$. We identify an initial state $s_{0}\in{\mathcal{S}}$ from which all the trajectories start. Moreover, since we are in a finite-horizon setting, all these trajectories are finite and we assume that they eventually end at an abstract terminal state $s_{f}\notin{\mathcal{S}}$ acting as a “sink” state. The state space is defined as a superset of the sample space ${\mathcal{X}}\subseteq{\mathcal{S}}$, and is structured in such a way that $x\in{\mathcal{X}}$ iff we can transition from $x$ to the terminal state $s_{f}$ (i.e., there exists an action $a\in{\mathcal{A}}$ such that $T(x,a)=s_{f}$); the states $x\in{\mathcal{X}}$ are called terminating states, following the naming convention of Bengio et al. (2023). We also set the discount factor $\gamma=1$ throughout this paper.
Since the MDP is deterministic, we can identify the action $a$ leading to a state $s^{\prime}=T(s,a)$ with the transition $s\rightarrow s^{\prime}$ in the state space itself. As such, we will write all quantities involving state-action pairs $(s,a)$ in terms of $(s,s^{\prime})$ instead. The reward function $r(s,s^{\prime})$ is defined such that the sum of rewards along a complete trajectory (the return) only depends on the energy of the terminating state it reaches: for a trajectory $(s_{0},s_{1},\ldots,s_{T},s_{f})$, we have
@@
with the convention $s_{T+1}=s_{f}$. In particular, this covers the case of a sparse reward that is received only at the end of the trajectory (i.e., $r(s_{T},s_{f})=-{\mathcal{E}}(s_{T})$, and zero everywhere else). This decomposition of the energy into intermediate rewards is similar to Buesing et al. (2020).
@@
$\pi^{*}_{\mathrm{MaxEnt}}=\operatorname*{arg\,max}_{\pi}\sum_{t=0}^{T}\mathbb{E}_{\tau}\big{[}r(s_{t},s_{t+1})+\alpha{\mathcal{H}}(\pi(\cdot\mid s_{t}))\big{]}.$ (3)
@@
In what follows, it will be convenient to view the transitions in this deterministic MDP as a directed acyclic graph (DAG) ${\mathcal{G}}$ over the states in ${\mathcal{S}}$ (including $s_{f}$), rooted in $s_{0}$; the graph is acyclic to ensure that ${\mathcal{M}}_{\mathrm{soft}}$ is finite-horizon. We will use the notations $\mathrm{Pa}(s)$ and $\mathrm{Ch}(s)$ to denote respectively the parents and the children of a state $s$ in ${\mathcal{G}}$. For any terminating state $x\in{\mathcal{X}}$, $s_{0}\rightsquigarrow x$ denotes a complete trajectory in ${\mathcal{G}}$ of the form $\tau=(s_{0},s_{1},\ldots,x,s_{f})$; the transition to the terminal state is implicit in the notation, albeit necessary in our definition.
@@ 2.2 Sampling terminating states from a soft MDP
From the literature on control as inference (Ziebart et al., 2008; Levine, 2018), it can be shown that the policy maximizing the MaxEnt RL objective in Equation 3 induces a distribution over dynamically consistent trajectories $\tau=(s_{0},s_{1},\ldots,s_{T},s_{f})$ that depends on the sum of rewards along the trajectory:
$\pi^{*}(\tau)\triangleq\prod_{t=0}^{T}\pi^{*}_{\mathrm{MaxEnt}}(s_{t+1}\mid s_{t})\propto\exp\left(\frac{1}{\alpha}\sum_{t=0}^{T}r(s_{t},s_{t+1})\right).$ (4)
With our choice of reward function in Equation 2, this suggests a simple strategy to sample from the Gibbs distribution in Equation 1, once the optimal policy $\pi^{*}_{\mathrm{MaxEnt}}$ is known: sample a trajectory in ${\mathcal{M}}_{\mathrm{soft}}$, starting at the initial state $s_{0}$ and following the optimal policy, and only return the terminating state $s_{T}$ reached at the end of $\tau$ (ignoring the terminal state $s_{f}$). However, in general, we are interested in sampling an object $x\in{\mathcal{X}}$ (the terminating state) using this sequential process, but not how this object is generated (the exact trajectory taken). The distribution of interest is therefore not a distribution over trajectories as in Equation 4, but its marginal over terminating states:
@@
$\pi^{*}(x)$ is called the terminating state distribution associated with the policy $\pi^{*}_{\mathrm{MaxEnt}}$ (Bengio et al., 2023). When the state transition graph ${\mathcal{G}}$ of the soft MDP is a tree [^1] With the exception of the terminal state $s_{f}$, whose parents are always all the terminating states in ${\mathcal{X}}$. rooted in $s_{0}$, and there is a unique complete trajectory $s_{0}\rightsquigarrow x$ leading to any $x\in{\mathcal{X}}$, then this process is guaranteed to be equivalent to sampling from the Gibbs distribution (Bengio et al., 2021). Examples of tree-structured MDPs include the autoregressive generation of sequences (Bachman and Precup, 2015; Weber et al., 2015; Angermueller et al., 2020; Jain et al., 2022; Feng et al., 2022), and sampling from a discrete factor graph with a fixed ordering of the random variables (Buesing et al., 2020).
Figure 1: Illustration of the bias of the terminating state distribution associated with $\pi^{*}_{\mathrm{MaxEnt}}$ on a soft MDP with a DAG structure. The labels on each transition of the MDP corresponds to the reward function, satisfying Equation 2 (sparse reward setting). The terminating state distribution $\pi^{*}(x)$ is computed by marginalizing $\pi^{*}(\tau)$ over trajectories leading to $x$ (e.g., two trajectories $s_{0}\rightarrow s_{1}\rightarrow x_{4}$ and $s_{0}\rightarrow s_{2}\rightarrow x_{4}$ to $x_{4}$). $\pi^{*}(\tau)$ is computed based on Equation 4, and we assume $\alpha=1$. The terminating state distribution $\pi^{*}(x)$ should be contrasted with the (target) Gibbs distribution $P(x)\propto\exp(-{\mathcal{E}}(x))$. The normalization constant is $Z^{\prime}=\exp(-{\mathcal{E}}(x_{3}))+2\exp(-{\mathcal{E}}(x_{4}))+\exp(-{\mathcal{E}}(x_{5}))$. This MDP is inspired by Jain et al. (2023a).
The equivalence between control and inference, or sampling, no longer holds when ${\mathcal{G}}$ is a general DAG. As an illustrative example, shown in Figure 1, consider the problem of generating small molecules by adding fragments one at a time, as described in You et al. (2018). If we follow the optimal policy, even though the distribution over trajectories is proportional to $\exp(-{\mathcal{E}}(x))$ as in Equation 4, its marginal over terminating states is biased since there are two trajectories leading to the same molecule $x_{4}$ (i.e., multiple orders in which fragments can be added). This bias was first highlighted by Bengio et al. (2021).
@@ 2.3 Generative Flow Networks
To address this mismatch between the target Gibbs distribution in Equation 1 and the terminating state distribution in Equation 5 induced by the optimal policy in MaxEnt RL, Bengio et al. (2021) introduced a new class of probabilistic models over discrete and compositional objects called Generative Flow Networks (GFlowNets; Bengio et al., 2023). Instead of searching for a policy maximizing Equation 3, the goal of a GFlowNet is to find a flow function $F(s\rightarrow s^{\prime})$ defined over the edges of ${\mathcal{G}}$ that satisfies the following flow-matching conditions for all states $s^{\prime}\in{\mathcal{S}}$ such that $s^{\prime}\neq s_{0}$:
@@
with an additional boundary condition $F(x\rightarrow s_{f})=\exp(-{\mathcal{E}}(x)/\alpha)$ for each terminating state $x\in{\mathcal{X}}$. Putting it in words, Equation 6 means that the total amount of flow going into any state $s^{\prime}$ has to be equal to the amount of flow going out of it, except for the initial state $s_{0}$ which acts as a single “source” for all the flow. We can define a policy [^2] Following the conventions from the GFlowNet literature, we will use the notation $P_{F}$ for this policy, also called the forward transition probability (Bengio et al., 2023), to distinguish it from the optimal policy $\pi^{*}_{\mathrm{MaxEnt}}$ in MaxEnt RL. from a flow function by simply normalizing the flows going out of any state $s_{t}$:
@@
Bengio et al. (2021) showed that if there exists a flow function satisfying the flow-matching conditions in Equation 6 as well as the boundary conditions, then sampling terminating states by following $P_{F}$ in the soft MDP as described in Section 2.2 is equivalent to sampling from the Gibbs distribution in Equation 1. In that case, the terminating state distribution associated with the policy $P_{F}$ satisfies $P_{F}(x)\propto\exp(-{\mathcal{E}}(x)/\alpha)$, for any terminating state $x\in{\mathcal{X}}$, and is not biased by the number of trajectory leading to $x$, contrary to the optimal policy in MaxEnt RL.

### v2

@@ 1 Introduction
Approximate probabilistic inference has seen a tremendous amount of progress, notably from a variational perspective coupled with deep neural networks. This is particularly true for continuous sample spaces, where the Evidence Lower Bound (ELBO) can be maximized with gradient methods, thanks to methods such as pathwise gradient estimation [Kingma and Welling, 2014, Rezende et al., 2014, Rezende and Mohamed, 2015]. In the case of discrete and highly structured sample spaces though, this “reparametrization trick” becomes more challenging since it often requires continuous relaxations of discrete distributions [Jang et al., 2017, Maddison et al., 2017, Mena et al., 2018]. In those cases, variational inference can also be carried out more generally thanks to score function estimation [Williams, 1992], albeit at the expense of high variance [Mohamed et al., 2020]. An alternative for approximate inference is through sampling methods, based on Markov chain Monte Carlo (MCMC; Hastings, 1970, Gelfand and Smith, 1990), when the target distribution is defined up to an intractable normalization constant.
@@
Bengio et al. [2021] introduced a new class of probabilistic models called Generative Flow Networks (GFlowNets), to approximate an unnormalized target distribution over discrete and structured sample spaces from a variational perspective [Malkin et al., 2023, Zimmermann et al., 2022]. GFlowNets treat sampling as a sequential decision making problem, heavily inspired by the literature in Reinforcement Learning (RL). Unlike RL though, which seeks an optimal policy maximizing the cumulative reward, the objective of a GFlowNet is to find a policy such that objects can be sampled proportionally to their cumulative reward. Nevertheless, this relationship led to a number of best practices from the RL literature being transferred into GFlowNets, such as the use of a replay buffer [Shen et al., 2023, Vemgal et al., 2023] and target network [Deleu et al., 2022], and advanced exploration strategies [Rector-Brooks et al., 2023].
@@
Although Bengio et al. [2021] proved that GFlowNets were exactly equivalent to maximum entropy RL (MaxEnt RL; Ziebart, 2010) in some specific cases which we will recall in Section 2, it has long been thought that this connection was only superficial in general. However recently, Tiapkin et al. [2024] showed that GFlowNets and MaxEnt RL are in fact one and the same, up to a correction of the reward function. This, along with other recent works [Mohammadpour et al., 2024, Niu et al., 2024], paved the way to show deeper connections between GFlowNets and MaxEnt RL algorithms. In this work, we extend this correction of the reward to a more general case, and establish novel equivalences between GFlowNet & MaxEnt RL objectives, notably between the widely used Trajectory Balance loss in the GFlowNet literature [Malkin et al., 2022], and the Path Consistency Learning algorithm in MaxEnt RL [Nachum et al., 2017]. We also introduce a variant of the Soft Q-Learning algorithm [Haarnoja et al., 2017], depending directly on a policy, and show that it becomes equivalent to the Modified Detailed Balance loss of Deleu et al. [2022]. Finally, we show these similarities in behavior empirically on three different domains, and include the popular Soft Actor-Critic algorithm (SAC; Haarnoja et al., 2018a) in our evaluations, which has no existing GFlowNet counterpart, under similar conditions on the same domains.
@@ 2.1 Maximum entropy Reinforcement Learning
We consider a finite-horizon Markov Decision Process (MDP) ${\mathcal{M}}_{\mathrm{soft}}=({\mathcal{S}},{\mathcal{A}},s_{0},T,r)$, where the state space ${\mathcal{S}}$ and the action space ${\mathcal{A}}$ are discrete and finite. We assume that this MDP is deterministic, with a transition function $T:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\bar{{\mathcal{S}}}$ that determines how to move to a new state $s^{\prime}=T(s,a)$ from the state $s$, following the action $a$. We identify an initial state $s_{0}\in{\mathcal{S}}$ from which all the trajectories start. Moreover, since we are in a finite-horizon setting, all these trajectories are finite and we assume that they eventually end at an abstract terminal state $s_{f}\notin{\mathcal{S}}$ acting as a “sink” state; we use the notation $\bar{{\mathcal{S}}}={\mathcal{S}}\cup\{s_{f}\}$. The state space is defined as a superset of the sample space ${\mathcal{X}}\subseteq{\mathcal{S}}$, and is structured in such a way that $x\in{\mathcal{X}}$ iff we can transition from $x$ to the terminal state $s_{f}$ (i.e., there exists an action $a\in{\mathcal{A}}$ such that $T(x,a)=s_{f}$); the states $x\in{\mathcal{X}}$ are called terminating states, following the naming convention of Bengio et al. [2023]. We also set the discount factor $\gamma=1$ throughout this paper. This setting with known deterministic dynamics is well-studied in the Reinforcement Learning literature [Todorov, 2006, Kappen et al., 2012].
Since the MDP is deterministic, we can identify the action $a$ leading to a state $s^{\prime}=T(s,a)$ with the transition $s\rightarrow s^{\prime}$ in the state space itself. As such, we will write all quantities involving state-action pairs $(s,a)$ in terms of $(s,s^{\prime})$ instead. The reward function $r(s,s^{\prime})$ is defined such that the sum of rewards along a complete trajectory (the return) only depends on the energy of the terminating state it reaches: for a trajectory $\tau=(s_{0},s_{1},\ldots,s_{T},s_{f})$, we have
@@
with the convention $s_{T+1}=s_{f}$. In particular, this covers the case of a sparse reward that is received only at the end of the trajectory (i.e., $r(s_{T},s_{f})=-{\mathcal{E}}(s_{T})$, and zero everywhere else). This decomposition of the energy into intermediate rewards is similar to Buesing et al. [2020].
@@
$\pi^{*}_{\mathrm{MaxEnt}}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\tau}\Bigg{[}\sum_{t=0}^{T}r(s_{t},s_{t+1})+\alpha{\mathcal{H}}(\pi(\cdot\mid s_{t}))\Bigg{]},$
where the expectation is taken over complete trajectories $\tau$ sampled following the policy $\pi$. Intuitively, adding the entropy to the objective encourages stochasticity in the optimal policy and improves diversity. To highlight the difference with the standard objective in RL, ${\mathcal{M}}_{\mathrm{soft}}$ will be called soft MDP, following the nomenclature of Ziebart [2010].
@@
In what follows, it will be convenient to view the transitions in the soft MDP as a directed acyclic graph (DAG) ${\mathcal{G}}$ over the states in $\bar{{\mathcal{S}}}$ (including $s_{f}$), rooted in $s_{0}$; the graph is acyclic to ensure that ${\mathcal{M}}_{\mathrm{soft}}$ is finite-horizon. We will use the notations $\mathrm{Pa}(s)$ and $\mathrm{Ch}(s)$ to denote respectively the parents and the children of a state $s$ in ${\mathcal{G}}$. For any terminating state $x\in{\mathcal{X}}$, $s_{0}\rightsquigarrow x$ denotes a complete trajectory in ${\mathcal{G}}$ of the form $\tau=(s_{0},s_{1},\ldots,x,s_{f})$; the transition to the terminal state is implicit in the notation, albeit necessary in our definition.
@@ 2.2 Sampling terminating states from a soft MDP
From the literature on control as inference [Ziebart et al., 2008, Levine, 2018], it can be shown that the policy maximizing the MaxEnt RL objective in LABEL:eq:maxent-rl-problem induces a distribution over dynamically consistent trajectories $\tau=(s_{0},s_{1},\ldots,s_{T},s_{f})$ that depends on the sum of rewards along the trajectory:
$\pi^{*}(\tau)\triangleq\prod_{t=0}^{T}\pi^{*}_{\mathrm{MaxEnt}}(s_{t+1}\mid s_{t})\propto\exp\left(\frac{1}{\alpha}\sum_{t=0}^{T}r(s_{t},s_{t+1})\right).$
With our choice of reward function in Equation 2, this suggests a simple strategy to sample from the Gibbs distribution in Equation 1, once the optimal policy $\pi^{*}_{\mathrm{MaxEnt}}$ is known: sample a trajectory in ${\mathcal{M}}_{\mathrm{soft}}$, starting at the initial state $s_{0}$ and following the optimal policy, and only return the terminating state $s_{T}$ reached at the end of $\tau$ (ignoring the terminal state $s_{f}$). However, in general, we are interested in sampling an object $x\in{\mathcal{X}}$ (the terminating state) using this sequential process, but not how this object is generated (the exact trajectory taken). The distribution of interest is therefore not a distribution over trajectories as in LABEL:eq:distribution-trajectories, but its marginal over terminating states:
@@
$\pi^{*}(x)$ is called the terminating state distribution associated with the policy $\pi^{*}_{\mathrm{MaxEnt}}$ [Bengio et al., 2023]. When the state transition graph ${\mathcal{G}}$ of the soft MDP is a tree [^1] With the exception of the terminal state $s_{f}$, whose parents are always all the terminating states in ${\mathcal{X}}$. rooted in $s_{0}$, and there is a unique complete trajectory $s_{0}\rightsquigarrow x$ leading to any $x\in{\mathcal{X}}$, then this process is guaranteed to be equivalent to sampling from the Gibbs distribution [Bengio et al., 2021]. Examples of tree-structured MDPs include the autoregressive generation of sequences [Bachman and Precup, 2015, Weber et al., 2015, Angermueller et al., 2020, Jain et al., 2022, Feng et al., 2022], and sampling from a discrete factor graph with a fixed ordering of the random variables [Buesing et al., 2020].
Figure 1: Illustration of the bias of the terminating state distribution associated with $\pi^{*}_{\mathrm{MaxEnt}}$ on a soft MDP with a DAG structure. The labels on each transition of the MDP corresponds to the reward function, satisfying Equation 2 (sparse reward setting). The terminating state distribution $\pi^{*}(x)$ is computed by marginalizing $\pi^{*}(\tau)$ over trajectories leading to $x$ (e.g., two trajectories $s_{0}\rightarrow s_{1}\rightarrow x_{4}$ and $s_{0}\rightarrow s_{2}\rightarrow x_{4}$ to $x_{4}$). $\pi^{*}(\tau)$ is computed based on LABEL:eq:distribution-trajectories, and we assume $\alpha=1$. The terminating state distribution $\pi^{*}(x)$ should be contrasted with the (target) Gibbs distribution $P(x)\propto\exp(-{\mathcal{E}}(x))$. The normalization constant is $Z^{\prime}=\exp(-{\mathcal{E}}(x_{3}))+2\exp(-{\mathcal{E}}(x_{4}))+\exp(-{\mathcal{E}}(x_{5}))$. This MDP is inspired by Jain et al. [2023a].
The equivalence between control and inference (sampling from Equation 1) no longer holds when ${\mathcal{G}}$ is a general DAG though. As an illustrative example, shown in Figure 1, consider the problem of generating small molecules by adding fragments one at a time, as described in You et al. [2018]. If we follow the optimal policy, even though the distribution over trajectories is proportional to $\exp(-{\mathcal{E}}(x))$ as in LABEL:eq:distribution-trajectories, its marginal over terminating states is biased since there are two trajectories leading to the same molecule $x_{4}$ (i.e., multiple orders in which fragments can be added). This bias was first highlighted by Bengio et al. [2021].
@@ 2.3 Generative Flow Networks
To address this mismatch between the target Gibbs distribution in Equation 1 and the terminating state distribution in Equation 5 induced by the optimal policy in MaxEnt RL, Bengio et al. [2021] introduced a new class of probabilistic models over discrete and compositional objects called Generative Flow Networks (GFlowNets; Bengio et al., 2023). Instead of searching for a policy maximizing LABEL:eq:maxent-rl-problem, the goal of a GFlowNet is to find a flow function $F(s\rightarrow s^{\prime})$ defined over the edges of ${\mathcal{G}}$ that satisfies the following flow-matching conditions for all states $s^{\prime}\in{\mathcal{S}}$ such that $s^{\prime}\neq s_{0}$:
@@
with an additional boundary condition for each terminating state $x\in{\mathcal{X}}$: $F(x\rightarrow s_{f})=\exp(-{\mathcal{E}}(x)/\alpha)$. Putting it in words, Equation 6 means that the total amount of flow going into any state $s^{\prime}$ has to be equal to the amount of flow going out of it, except for the initial state $s_{0}$ which acts as a single “source” for all the flow. We can define a policy [^2] Following the conventions from the GFlowNet literature, we will use the notation $P_{F}$ for this policy, also called the forward transition probability [Bengio et al., 2023], to distinguish it from the optimal policy $\pi^{*}_{\mathrm{MaxEnt}}$ in MaxEnt RL. from a flow function by simply normalizing the flows going out of any state $s_{t}$:
@@
Bengio et al. [2021] showed that if there exists a flow function satisfying the flow-matching conditions in Equation 6 as well as the boundary conditions, then sampling terminating states by following $P_{F}$ in the soft MDP as described in Section 2.2 is equivalent to sampling from Equation 1. In that case, the terminating state distribution associated with the policy $P_{F}$ satisfies $P_{F}(x)\propto\exp(-{\mathcal{E}}(x)/\alpha)$, for any terminating state $x\in{\mathcal{X}}$, and is not biased by the number of trajectory leading to $x$, contrary to the optimal policy in MaxEnt RL.
[…]
Figure 2: Equivalence between objectives in MaxEnt RL, with corrected rewards, and the objectives in GFlowNets. The objectives are classified based on whether they operate at the level of (complete) trajectories (left), transitions (middle), or if all the states are terminating (right). Further details about the form of the different residuals and the correspondences to transfer from one objective to another are available in Figure 6.